Find three consecutive odd integers such that the sum of the first, two times the second, and three times the third is 70.
Answer by mananth(16946)

let the integers be x, x+2, and x+4
x+2(x+2)+3(x+4)=70
x+2x+4+3x+12=70
6x+16=70
6x=70-16
6x=54
x=9
the numbers are 9,11,13